# Predictability of Performance in Communication Networks Under Markovian Dynamics

### v3

@@ Abstract
With the emergence of time-critical applications in modern communication networks, there is a growing demand for proactive network adaptation and quality of service (QoS) prediction. However, a fundamental question remains largely unexplored: how can we quantify and achieve more predictable communication systems in terms of performance? To address this gap, this paper introduces a theoretical framework for defining and analyzing predictability in communication systems, with a focus on the impact of observations for performance forecasting. We establish a mathematical definition of predictability based on the total variation distance between forecast and marginal performance distributions. A system is deemed unpredictable when the forecast distribution, providing the most comprehensive characterization of future states using all accessible information, is indistinguishable from the marginal distribution, which depicts the system’s behavior without any observational input. This framework is applied to multi-hop systems under Markovian conditions, with a detailed analysis of Geo/Geo/1 queuing models in both single-hop and multi-hop scenarios. We derive exact and approximate expressions for predictability in these systems, as well as upper bounds based on spectral analysis of the underlying Markov chains. Our results have implications for the design of efficient monitoring and prediction mechanisms in future communication networks aiming to provide dependable services.
@@ I Introduction
With the rapid growth of the number of mobile communication users and applications in recent decades, the capacity and data rates of wireless networks have increased significantly. This progress has been achieved incrementally through successive generations (e.g., from 3G to 5G), which have substantially boosted achievable throughput. In more recent years, the demand for time-critical applications has emerged with stringent requirements regarding the end-to-end delay of the communication system [1].
These applications, usually categorized into cyber-physical systems and human-in-the-loop applications, rely on the network’s end-to-end quality of service (QoS). In cyber-physical systems, scenarios such as vehicle to everything (V2X), industrial internet of things (IIoT), etc are considered where large packet delays may trigger system instability and jeopardize the safety of the system. In human-in-the-loop applications, such as those involving extended reality (XR) headsets, users demand a high level of responsiveness, and substantial packet delays can lead to motion sickness [2].
In response to these challenges, the community has established quality of service (QoS) criteria tailored to individual applications, which is standardized by 3GPP in 5G and continues to be refined for 6G. These criteria are defined by setting a delay target and a reliability level that limits the delay violation probability (DVP). For instance, a delay target of 10 ms with a reliability of 0.9999 has been set for industrial control applications [3, 4].
The stringent requirements of the majority of these use cases ask for proactive network adaptation approaches via predicting QoS KPIs such as end-to-end delay in advance. Proactive adaptation can offer major gains to the operation of these applications, for instance, predicting in advance a potential increase in the network delay and/or deterioration of computing resources will enable the network and/or the respective application to adapt and achieve the required quality of service [5, 6, 7, 8].
@@
Consequently, there is a strong focus in the development of 6G on utilizing end-to-end delay probability predictors, especially in the integration with time sensitive networking (TSN) [9, 10]. Moreover, 5G Automotive Association (5GAA) has introduced the concept of predictive QoS, which refers to the mechanisms enabling mobile networks to provide notifications about predicted QoS changes to interested consumers in advance [11]. ETSI, in the context of mobile edge computing (MEC), has also introduced the notion of predictive QoS support [12]. In this context, the prediction of potential handovers leading to the estimated QoS performance is described as the key solution for mitigating any service interruption and support the application requirements.
@@
Predictability in this context refers to the ability to anticipate the future performance of the system based on current and historical data. As the communication system interacts with the environment, the predictability of the system extends beyond just the immediate network parameters. If the environmental conditions—such as interference patterns, user mobility, or traffic loads—exhibit non-random behavior and can be forecasted with reasonable accuracy, this adds a new dimension to the system’s overall predictability. Therefore, it is essential to understand how a communication system can be considered predictable, and to identify the factors that enhance or undermine this predictability.
@@
  • The predictability derivations are applied to an abstract queuing system, focusing on a Geo/Geo/1 queue in both single and multi-hop scenarios. In this analysis of end-to-end delay, queue length observations are considered the primary predictive factor. The study includes exact and approximate derivations of predictability.
  • Finally, by numerical evaluations, we showcased the utility of our predictability analysis in the queuing setup. The evaluations investigate the trade-offs between predictability and observability in single and multi-hop queueing systems. We demonstrated the accuracy of our proposed predictability measures and approximations by comparing them with exact calculations.
@@ II Related Works
At the most basic level, there are works that define predictability in a setup where point prediction of a univariate metric is desired. Autocorrelation which evaluates how well a time series correlates with its own time-lagged versions, is a widely used indicator of predictability in finance [13]. Permutation entropy (PE) is the metric employed in fields like ecology and physics, to gauge predictability [14, 15, 16]. PE measures time series complexity by quantifying the diversity and frequency of ordinal patterns formed by subsequences of the series. A higher permutation entropy indicates more diverse motifs, leading to lower predictability. This indicator is mostly empirically evaluated and has been applied to various domains, including predicting infectious disease outbreaks [17], and ecological systems [16]. Abeliuk et al. explore the relationship between predictability and sampling in partially observed systems. They evaluate multiple predictability measures such as PE and autocorrelation in outbreaks of infectious diseases, digital forum interactions, and software development projects [18].
@@
Besides autocorrelation and PE, another common approach in defining predictability is to find the lower bound on prediction error [19]. In some works, the empirically determined entropy of the target metric is incorporated in Fano’s inequality to derive the bound on error. Using this approach Li et al. investigate the limits of predictability for urban vehicular location and staying time [20]. More information-theoretic measures of predictability have inspired numerous studies. For instance, the studies on the predictability of human mobility and conversation use random entropy, uncorrelated entropy, and conditional entropy to measure predictability [21, 22].
@@
More advanced studies in this category incorporate side information or covariates into the predictability model. As a result, the selected predictability measure is typically relative entropy, conditional entropy, or mutual information between the side information and the target metric. Bialek et al. define predictive information as the mutual information between past and future states of a time series as a criterion for predictability [23]. Haven et al. used a Gaussian ensemble prediction to assess relative entropy for predictability analysis [24]. More recently, Li et al. applied conditional entropy to traffic forecasting, highlighting the limitations of traffic predictability [25]. Also, Fang and Lee explore the predictability in the context of regression problems in machine learning. They analyze predictability in this context using the conditional entropy between the features and the label and introduce lower and upper bounds for it [26]. Due to the incorporation of side information, these predictability analysis are considered more comprehensive in our perspective.
@@
In atmospheric sciences literature, where the probabilistic forecasting of the weather systems is of interest, the information-theoretic definition of predictability with side information is gaining more attention [27, 28]. In a series of papers, DelSole et al. define predictability as the difference between forecast and climatological distributions (marginal distribution), with the latter serving as the baseline for probabilistic information about the metric where no side information is incorporated. Various information theory measures were proposed and analyzed to quantify predictability, in systems with linear stochastic models and Gaussian processes. [29].
Among the works that are related to networking and communication systems, Ding et al. examine the predictability of radio spectrum state dynamics using Fano’s inequality. Their findings show high predictability in real-world spectrum measurements, which has significant implications for cognitive radio networks and 5G spectrum sharing [30]. In another paper, Sang et al. apply predictability analysis to an abstract network model with a queue. They investigate the predictability of network traffic by analyzing how far into the future traffic rates can be predicted with a given error constraint on the prediction problem [31].
@@
Despite the growing focus on quality of service prediction in communication systems, comprehensive predictability analyses are both necessary and, to the best of our knowledge, largely nonexistent. This work addresses this gap by adopting a thorough predictability model and demonstrating its implications for communication system performance. Drawing inspiration from the predictability framework used in atmospheric science, which is an information-theoretic measure well-suited for probabilistic forecasting, our definition compares the conditional distribution of performance given the side information or observations with its marginal distribution. Such definition offers great capabilities not only in applying it to various dynamical systems, but also we show how it can model imperfect observability. Moreover, since Markov chains are widely used in queue modeling and various fields—including stochastic processes and machine learning—due to their effectiveness in modeling dynamic systems with memoryless state transitions, we employ the Markovian assumption for the system’s dynamics. In the following section we step into system model and predictability definition in more detail.
@@ III System Model and Problem Formulation
Figure 1: Multi-hop communication system model with observable measures being conditions and performance.
Forecasting performance in a communication network becomes more manageable and precise by breaking down the system into independent modules and modeling each individually. This also allows for more efficient proactive adjustments, as we can more accurately pinpoint the sources of performance variations or declines [32]. These components can correspond to tandem data transmission links or, within a single link, to multiple processes that sequentially handle data bits to ensure reliable and efficient communication. Furthermore, the condition or state of each subsystem can be monitored and, according to domain knowledge, correlated with performance.
@@
We consider a communication system that consists of $M$ independent subsystems or hops as shown in Figure 1. In time slot $n$ (time is slotted), subsystem $m$ exhibits a performance measure denoted by $Z^{(m)}_{n}$ which is a random variable over the discrete space $\mathcal{Z}$, for example, packet delay. $\Pr(Z^{(m)}_{n})$ denotes the marginal distribution (or prior distribution) and is considered stationary. It constitutes a baseline for the information content about performance, disregarding all observations. This distribution can be obtained by random sampling of the performance. In case of end-to-end delay (without packet drops) total performance is obtained via $Z_{n}=\sum_{m=1}^{M}Z_{n}^{(m)}$.
Moreover, each subsystem has an internal state or condition measure denoted by $X^{(m)}_{n}$ that influences performance, for example, the congestion level in the subsystem. Conditions are defined over a discrete finite state space $\mathcal{X}^{(m)}$ of size $K^{(m)}$. Then, the collection of all the state variables belonging to each subsystem is denoted by $X_{n}=\{X^{(1)}_{n},...,X^{(m)}_{n},...,X^{(M)}_{n}\}$ with size $K$. Observations of state measures across the system collected up to time $n$ are denoted by $O_{0:n}$. Depending on the scenario, observations can include all, parts, or an imperfect version of the conditions.
@@
We consider three different observability defects in our model: first, delayed observations, which is particularly crucial when a network of remote nodes is involved in the packet route. In this scenario, monitoring information would arrive with a delay to the prediction and adaptation point, i.e., $O_{n}=X_{n-d}$. In the second case, we may choose to ignore certain subsystems and not observe their state, i.e., $O_{n}\subset X_{n}$ due to the overhead that the monitoring task introduces. The third case, which also concerns overhead, mandates aggregated states. In this scenario, we assume states of $\mathcal{X}$ are aggregated to $\mathcal{A}$, i.e., there exists a surjective map $\phi:\mathcal{X}\rightarrow\mathcal{A}$ with $|\mathcal{A}|\leq|\mathcal{X}|$ (or $\bar{K}\leq K$ as $|\mathcal{A}|=\bar{K}$). Then, the observations are coduncted on the aggregated states $O_{n}=\phi(X_{n})$.
The distribution of performance right after observations become available denoted by $\Pr(Z_{n}\mid O_{0:n})$ is the posterior distribution [33]. The marginal distribution is related to the posterior distribution by the chain rule as it is obtained by integrating out all observations:
The primary objective of any performance predictor in this setting is to capture the relationship between the observations and future performance. The probability distribution of the performance metric $L$ timeslots into the future, based on the observations collected up to the present moment, is the forecast distribution:
$\Pr(Z_{n+L}\mid O_{0:n}=o_{0:n}).$ (2)
We examine performance in communication systems that their conditions adhere to the Markovian property, which is applicable in numerous contexts and mathematically manageable. We define two sets of conditional independence assumptions on the system: first, $Z_{n}$ is independent of all other condition states except the current one $X_{n}$ and second, $X_{n}$ is independent of $X_{1},...,X_{n-2}$ given $X_{n-1}$ (the first-order Markov property), i.e.,

### v4

@@ Abstract
With the emergence of time-critical applications in modern communication networks such as vehicle-to-everything (V2X) systems, there is a growing demand for proactive network adaptation and quality of service (QoS) prediction. However, a fundamental question remains largely unexplored: How can we quantify and achieve more predictable communication systems in terms of performance? To address this gap, this paper introduces a theoretical framework for defining and analyzing predictability in communication systems, with a focus on the impact of observations for performance forecasting. We establish a mathematical definition of predictability based on the total variation distance between the forecast and marginal performance distributions. A system is deemed unpredictable when the forecast distribution, providing the most comprehensive characterization of future states using all accessible information, is indistinguishable from the marginal distribution, which depicts the system’s behavior without any observational input. This framework is applied to multi-hop systems under Markovian conditions, with a detailed analysis of Geo/Geo/1/K queuing models in both single-hop and multi-hop scenarios. Additionally, we apply the framework to a random-walk-based model of QoS for connected vehicles experiencing changing channel conditions. We derive exact and approximate expressions for predictability in these systems, as well as upper bounds based on spectral analysis of the underlying Markov chains. Our results have implications for the design of efficient monitoring and prediction mechanisms in future communication networks aiming to provide dependable services.
@@ I Introduction
Over recent decades, the rapid expansion of mobile communication users and applications has intensified the demand for mission-critical services in 5G and advanced wireless networks. These applications—from industrial automation and remote surgery to extended reality (XR) —impose stringent performance requirements, particularly regarding end-to-end latency and reliability [1, 2].
To accommodate the diverse needs of these services, the community has developed tailored QoS criteria. Standardized by 3GPP in 5G and continually refined for 6G, these criteria set explicit delay targets and reliability levels that limit the delay violation probability (DVP). For example, industrial control applications typically require a delay target of 10 ms and a reliability of 99.99% [3, 4]. In V2X scenarios—encompassing cooperative driving, collision avoidance, and autonomous navigation—networks must support latencies ranging from sub-millisecond to a few milliseconds while achieving reliability levels of 99.999% or higher [5].
Meeting these stringent requirements is particularly challenging in networks characterized by high mobility and rapidly changing conditions, such as those encountered in V2X communications. Even occasional increases in delay can lead to significant disruptions or safety hazards. This challenge highlights both the need and the opportunity to develop accurate, proactive QoS prediction mechanisms capable of forecasting network performance degradation, preventing accidents, and maintaining robust control loops [6, 7, 8, 9].
@@
Consequently, the 5G automotive association (5GAA) introduced the concept of predictive QoS, which refers to the mechanisms that enable mobile networks to provide notifications about predicted QoS changes to interested consumers in advance [10]. The prediction of potential handovers leading to the estimated QoS performance is the key solution for mitigating any service interruption and supporting the V2X communication requirements. ETSI, in the context of mobile edge computing (MEC), has also introduced the notion of predictive QoS support [11] and there is a strong focus in the development of 6G on utilizing network latency predictors, especially in the integration with time sensitive networking (TSN), to meet the stringent requirements in scenarios such as V2X [12, 13].
A substantial body of research has explored performance prediction for V2X systems using machine learning techniques, examining a wide range of methods and applications [14, 15, 16, 17, 18, 19]. In many studies, network-related metrics such as signal to interference and noise ratio (SINR), channel quality indicator (CQI), and network queue state served as critical inputs to throughput or delay predictive models. Additionally, the notion of Age of Information (AoI), which quantifies how fresh or timely the most recently received data are, can also play a role in maintaining prediction accuracy, as outdated information may limit how accurately these models can forecast system performance [20]. Also, including dynamic factors like speed and location of the user equipment (UE) or vehicle has proven essential for improving accuracy. However, we contend that the confidence of performance prediction models is inherently limited by the accuracy of their inputs. For instance, when considering dynamic factors, high UE mobility is expected to cause rapid fluctuations in signal strength and increased variability in network delays, whereas low mobility generally results in more stable delays. This can also be analyzed in the context of network-related metrics such as queue lengths; understanding how they evolve over time, improves our ability to assess the predictability of the performance of the network.
@@
Predictability in this context refers to the ability to anticipate the system’s future performance based on current and historical data. As the communication system interacts with the environment, the predictability of the system extends beyond just the immediate network parameters. Suppose the environmental conditions—such as interference patterns, user mobility, or traffic loads—exhibit non-random behavior and can be forecasted with reasonable accuracy. In that case, this adds a new dimension to the system’s overall predictability. Therefore, it is essential to understand how a communication system can be considered predictable and to identify the factors that enhance or undermine this predictability.
@@
  • We apply our predictability derivations to a simplified queueing model, focusing on a Geo/Geo/1/K system in both single-hop and multi-hop settings. The primary performance metric is the end-to-end delay, and we also address packet loss caused by blocking. In each case, the queue length serves as the predictive feature of interest. The study includes exact and approximate derivations of predictability for this queuing system.
  • Deriving closed-form expressions for the total variation distance between forecast and marginal distributions in Markov-modulated processes poses significant mathematical challenges. In this work, we introduce a spectral-based upper bound on predictability that exploits the Markov chain’s mixing behavior, and we further develop an approximation method for the Geo/Geo/1/K queue. These techniques reduce computational overhead while retaining analytical clarity, thereby facilitating tractable performance and predictability studies in a broad range of networking scenarios.
  • To extend this analysis to vehicular networks, we apply our framework to evaluate the predictability of network throughput under a random walk model of CQI, capturing how connected vehicles experience changing channel conditions.
  • Finally, by numerical evaluations, we showcased the utility of our predictability analysis in both queuing and random walk setup. The evaluations investigate the trade-offs between predictability and observability in single and multi-hop queueing systems. We demonstrated the accuracy of our proposed predictability measures and approximations by comparing them with exact calculations.
@@ II Related Works
At the most basic level, there are works that define predictability in a setup where point prediction of a univariate metric is desired. Autocorrelation which evaluates how well a time series correlates with its own time-lagged versions, is a widely used indicator of predictability in finance [21]. Permutation entropy (PE) is a metric employed in fields like ecology and physics to gauge predictability [22, 23, 24]. PE measures time series complexity by quantifying the diversity and frequency of ordinal patterns formed by subsequences of the series. A higher permutation entropy indicates more diverse motifs, leading to lower predictability. This indicator is mostly empirically evaluated and has been applied to various domains, including predicting infectious disease outbreaks [25], and ecological systems [24]. Abeliuk et al. explore the relationship between predictability and sampling in partially observed systems. They evaluate multiple predictability measures such as PE and autocorrelation in outbreaks of infectious diseases, digital forum interactions, and software development projects [26].
@@
Besides autocorrelation and PE, another common approach in defining predictability is to find the lower bound on prediction error. In some works, the empirically determined entropy of the target metric is incorporated in Fano’s inequality to derive the bound on error. Using this approach Li et al. investigate the limits of predictability for urban vehicular location and staying time [27]. More information-theoretic measures of predictability have inspired numerous studies. For instance, the studies on the predictability of human mobility and conversation use random entropy, uncorrelated entropy, and conditional entropy to measure predictability [28, 29].
@@
More advanced studies in this category incorporate side information or covariates into the predictability model. As a result, the selected predictability measure is typically relative entropy, conditional entropy, or mutual information between the side information and the target metric. Bialek et al. define predictive information as the mutual information between past and future states of a time series as a criterion for predictability [30]. Haven et al. use a Gaussian ensemble prediction to assess relative entropy for predictability analysis [31]. More recently, Li et al. apply conditional entropy to traffic forecasting, highlighting the limitations of traffic predictability [32]. Also, Fang and Lee explore the predictability in the context of regression problems in machine learning. They analyze predictability in this context using the conditional entropy between the features and the label and introduce lower and upper bounds for it [33]. Due to the incorporation of side information, these predictability analysis are considered more comprehensive in our perspective.
@@
In atmospheric sciences literature, where the probabilistic forecasting of the weather systems is of interest, the information-theoretic definition of predictability with side information is gaining more attention [34, 35]. In a series of papers, DelSole et al. define predictability as the difference between forecast and climatological distributions (marginal distribution), with the latter serving as the baseline for probabilistic information about the metric where no side information is incorporated. Various information-theoretic measures have been proposed and analyzed to quantify predictability in systems with linear stochastic models and Gaussian processes. [36].
Among the works that are related to networking and communication systems, Ding et al. examine the predictability of radio spectrum state dynamics using Fano’s inequality. Their findings show high predictability in real-world spectrum measurements, which has significant implications for cognitive radio networks and 5G spectrum sharing [37]. In another paper, Sihai et al. analyze voice traffic predictability in wireless networks using entropy, which captures temporal correlations in call detail records. Their study establishes a theoretical upper bound on the accuracy and demonstrates that using N-order Markov models [38]. Jing et al. study structural predictability in directed networks, linking it to controllability. They define predictability empirically through link prediction accuracy, using a leave-one-out ranking method. Unlike entropy-based approaches, their work offers no theoretical bounds but highlights the role of network structure in predictability [39].
@@
Despite the growing focus on quality of service prediction in communication systems, comprehensive predictability analyses are both necessary and, to the best of our knowledge, largely nonexistent. This work addresses this gap by adopting a thorough predictability model and demonstrating its implications for communication system performance. Drawing inspiration from the predictability definition used in atmospheric science, which is an information-theoretic framework well-suited for probabilistic forecasting, our definition compares the conditional distribution of performance given the side information or observations with its marginal distribution. Such definition offers great capabilities not only in applying it to various dynamical systems, but also we show how it can model imperfect observability.
For the choice of predictability measure, no single measure is universally optimal; each measure provides unique insights depending on the specific system and the goals of the analysis. Autocorrelation effectively captures linear dependencies in time-series data, while permutation entropy excels at quantifying non-linear structure in univariate signals. Entropy-based measures (e.g., cross entropy or entropy difference) provide a broad information-theoretic perspective—especially valuable when side information is involved—by quantifying the “surprise” or extra coding cost when using one distribution (forecast) in place of another (marginal). Entropy-based measures are highly sensitive to low-probability mismatches but are not metrics, as they generally lack symmetry and not satisfy the triangle inequality. In contrast, total variation (TV) distance is a true metric that directly measures the maximum difference between two distributions, remains bounded even if one distribution assigns negligible probability to events the other deems likely, and is subadditive. Subadditivity proves particularly useful in “systems of systems” (e.g., multi-hop communication networks) and Markovian settings: the overall discrepancy is no greater than the sum of discrepancies in each subsystem. Meanwhile, entropy-based measures offer complementary insights into average-case information loss, underscoring that the choice of measure should reflect whether one prioritizes worst-case system-level control (favoring TV distance) or average-case informational properties (favoring entropy-based approaches) [40].
@@ III System Model and Problem Formulation
Forecasting performance in a communication network becomes more manageable and precise by breaking down the system into independent modules and modeling each individually allows for more efficient proactive adjustments, as it enables more precise identification of performance variations or declines [41]. These components can correspond to tandem data transmission links or, within a single link, to multiple processes that sequentially handle data bits to ensure reliable and efficient communication. Furthermore, the condition or state of each subsystem can be monitored and, according to domain knowledge, correlated with performance. Before examining the system model and its predictability definition in detail, the following subsection provides a concise overview of the main parameters and their associated notations.
III-A Notations
In this subsection, we summarize the main notations and symbols used throughout the paper to facilitate clarity:
  • $\Pr(\cdot)$: Probability distribution. For instance, $\Pr(Z_{n})$ is the probability distribution of the random variable $Z_{n}$.
  • $\Pr(A\mid B)$: Conditional probability of event $A$ given event $B$.
  • $\lvert\mathcal{X}\rvert$: Cardinality of the set $\mathcal{X}$, i.e., the number of elements in $\mathcal{X}$.
  • $\|p-q\|_{\text{TV}}$: TV distance between distributions $p$ and $q$.
Moreover, Table I provides brief explanations of the key symbols used throughout the paper.
TABLE I: Key Model Parameters and Their Meanings
Symbol Meaning
$n$ Time slot index (discrete-time)
$m$, $M$ Subsystem index and number of subsystems (hops)
$Z_{n}$ / $Z_{n}^{(m)}$ Performance metric r.v. at time slot $n$ (subsystem $m$)
$\mathcal{Z}$ Space of performance r.v. $Z_{n}$
$X_{n}$ / $X_{n}^{(m)}$ Condition r.v. at time slot $n$ (subsystem $m$)
$\mathcal{X}$ / $\mathcal{X}^{(m)}$ Space of condition r.v. (size $K$ or $K^{(m)}$)
$\mathcal{A}$ / $\mathcal{A}^{(m)}$ Space of aggegated conditions r.v. (size $\bar{K}$ or $\bar{K}^{(m)}$)
$O_{n}$ Condition observation r.v. at time slot $n$
$O_{0:n}$ Collection of condition observations r.v. up to time $n$
$L$ Forecast lead time (number of slots ahead)
$d$ Observation delay (in slots)
$r_{y}(z)$ PMF of the posterior distribution
$P$ Markov chain’s transition probability matrix
$\pi$ Markov chain’s stationary distribution
$\alpha,\mu$ Arrival and service probabilities
$\rho=\alpha/\mu$ Utilization factor of the queue
@@ III-A Notations
Figure 1: Multi-hop communication system model with observable measures being conditions.
$X_{n-1}^{(m)}$ $X_{n}^{(m)}$ $X_{n+1}^{(m)}$ $Z_{n-1}^{(m)}$ $Z_{n}^{(m)}$ $Z_{n+1}^{(m)}$$\dots$
Figure 2: Temporal dependencies in the performance model for each subsystem $m$, underscoring the conditions’ Markov chain
III-B Multi-hop Communication System Model
@@
We consider a communication system that consists of $M$ independent subsystems or hops as shown in Figure 1. As packets traverse this multi-hop network, in time slot $n$ (time is slotted), subsystem $m$ exhibits a performance measure denoted by $Z^{(m)}_{n}$ which is a random variable over the discrete space $\mathcal{Z}$, for example, it can denote the delay of subsystem $m$. In this case, the end-to-end delay of the system (without packet drops) is obtained via $Z_{n}=\sum_{m=1}^{M}Z_{n}^{(m)}$. This multi-hop structure can, for instance, represent an autonomous vehicle scenario in which sensor data must first traverse a wireless uplink from the vehicle to a remote control server (hop 1), undergo processing to generate control commands (hop 2), and then return via a wireless downlink (hop 3) to the vehicle. Each hop introduces delay, and low end-to-end latency is crucial for guaranteeing system safety.
Moreover, each subsystem has an internal state or condition measure, denoted by $X^{(m)}_{n}$ which influences its delay. For example, the compute subsystem may experience congestion, while a wireless link hop might be subject to unfavorable fading conditions, leading to significant packet delays. These states lie in a discrete finite space $\mathcal{X}^{(m)}$ of size $K^{(m)}$. Collectively, all subsystem state variables form the overall system state, $X_{n}=\{X^{(1)}_{n},...,X^{(m)}_{n},...,X^{(M)}_{n}\}$ defined over a discrete space $\mathcal{X}$ with size $K$.
We assume a stationary stochastic dependency between the state and the performance of each subsystem, i.e.,
$\forall L,\Pr(Z_{n+L}=z|X_{n+L}=x)=\Pr(Z_{n}=z|X_{n}=x).$ (1)
This definition of stationarity means that the conditional distribution of performance given the state does not change over time, and it is fundamental to able to learn from past data and predict future performance. We argue that stationarity is a reasonable assumption that holds over extended periods of time. We denote the conditional distribution of the performance by $r_{x}(z)=\Pr(Z_{n}=z\mid X_{n}=x)$.
We consider that each subsystem’s state can be observed, and observations can be transmitted to a central server. This assumption is motivated by advancements in next-generation wireless technologies (e.g., 5G, 6G), which emphasize real-time monitoring of network conditions to enhance QoS [42]. We denote the observations of subsystem states up to time $n$ by $O_{0:n}$. Depending on the scenario, these observations may represent all, part, or an imperfect version of the true subsystem states. This assumption also recognizes that as monitoring capabilities grow, they inevitably introduce overhead into the system, requiring optimization of their resource usage. For example, SINR measurements and mobility data may be collected and uploaded via the wireless uplink to a monitoring server for decision-making. These monitoring links themselves can involve delays and lossy compression affecting accuracy.
@@
We consider three different observability defects in our model: first, delayed observations are particularly crucial when a network of wireless remote nodes is involved in the packet route. In this scenario, monitoring information would arrive with a delay to the prediction and adaptation point, i.e., $O_{n}=X_{n-d}$. In the second case, we may choose to ignore certain subsystems and not observe their state, i.e., $O_{n}\subset X_{n}$ due to the overhead that the monitoring task introduces. The third case, which also concerns overhead, mandates aggregated states. In this scenario, we assume states of $\mathcal{X}$ are aggregated to $\mathcal{A}$, i.e., there exists a surjective map $\phi:\mathcal{X}\rightarrow\mathcal{A}$ with $|\mathcal{A}|\leq|\mathcal{X}|$ (or $\bar{K}\leq K$ as $|\mathcal{A}|=\bar{K}$). Then, the observations are coduncted on the aggregated states $O_{n}=\phi(X_{n})$.
III-C Markovian Model of System State
We model the evolution of subsystem conditions over time as a first-order Markov process. This assumption posits that the future state of a subsystem, denoted by $X_{n}$, depends only on its current state, $X_{n-1}$, and the associated transition probabilities. Formally, this means that the probability of transitioning to a future state, $X_{n}=x_{i}$, given the entire history of states up to time $n-1$, is equivalent to the probability of transitioning to that state given only the current state:
$\Pr(X_{n}=x\mid X_{0},X_{1},\dots,X_{n-1})=\Pr(X_{n}=x\mid X_{n-1}).$ (2)
These transition probabilities, which govern the dynamics of the system’s state changes, encapsulate how likely a subsystem is to move from one condition to another in a single time step. The Markov property simplifies the analysis by abstracting the system’s history; the past only influences the future indirectly, through its contribution to the present state.
This Markovian model offers a tractable yet informative model for the evolution of network states in dynamic environments like vehicular and wireless networks. While real-world network behavior is complex, the Markov property allows us to focus on the most recent influences on a subsystem’s state and the typical patterns of change from that state. In high-mobility scenarios, the current location and surrounding environment of a UE often play a dominant role in determining short-term network states. Markovian models were used in works focusing on various aspects of network analysis and optimization, including vehicular network performance, channel allocation, and resource management [43, 44, 45]. While acknowledging the inherent limitations of this model, particularly regarding long-term dependencies, the Markov assumption provides a valuable balance between model accuracy and computational feasibility, enabling efficient analysis and real-time adaptation in dynamic network settings.
A direct result of the Markovian assumption and Equation 1 on the temporal dependencies in the performance model, as shown in Figure 2 is that